A proton and a deuteron with same initial kinetic energy enter a magnetic field in a direction perpendicular to the direction of the field.the ratio of the radii of the circular trajectories described by them is

Grade:12

2 Answers

Vikas TU
4 years ago
Hiii 
For Proton Let r(p) =x cm 
For deutron Let r(d)  = [2md(k)d]0.5/eB  = [4mp(k)p]0.50  = 1.41x
For alpha particles r(a) = [2ma(k)a]0.50/2eB = 2x cm 
So the ratio will be 1:sqrt2:2 or 1:1:sqrt2
Khimraj
4 years ago

Radius of trajectory is given by , r= MV / BQ

Similarly , V =

\sqrt{2vQ/M}

substitute and you will get ,

r= sqrt ( 2MV / Q)B ………………….. (1)

Here, V and B are constants … so we need to look into the mass and charge of each element.

  1. Proton mass is 1 a.m.u and charge is +1e
  2. For deutron , mass = 2 a.m.u
  3. Alpha partical has mass 4 a.m.u and charge 2e

Substituting these values in (1) you get the ratios as 1 : sqrt(2) : sqrt (2)